A simple pendulum is found to vibrate 50 times within 200seconds. When 1.5m of its length is reduced to a certain length, it vibrates 50 times in 175s. Find the original length of the pendulum.



1
Expert's answer
2022-01-23T16:15:22-0500

Solution

First case is

50 times within 200 s corresponds to a period of

"T=\frac{200}{50}=4sec."

The original length of the pendulum can be found from the period:

"T=2\pi \sqrt{\frac{l}{g}}"

So origional length

"l=\frac{gT^2}{4\pi^2}"

Putting all values

"l=\frac{9.8\times 4^2}{4\times(3.14)^2}"

"l=3.97m"